OpenStudy (anonymous):

Will Medal and fan! Thanks :) Solve. One point for showing the steps to solve each equation and one point for the correct solution. 4(x - 6) = -2(x - 3

OpenStudy (phi):

can you "distribute" the 4 on the left side? (that means multiply 4 times each term inside the parens)

OpenStudy (anonymous):

umm sure

OpenStudy (anonymous):

4*6 = 24 @Philospher

OpenStudy (anonymous):

oops sorry i tagged you

OpenStudy (anonymous):

I feel stupid...

OpenStudy (anonymous):

4x-24=-2x+6 then 6x=30 then x =5 and its done :D

OpenStudy (anonymous):

well, glad that i could help . Have a nice day :D

OpenStudy (anonymous):

Thank you!!